# How to Purchase Labels? A Cost-effective Approach Using Active Learning Markets

### v1

@@ 3.3 Data purchase strategies
The data purchase problem addressed in this paper is inherently constrained by budget and governed by heterogeneous pricing. Active learning, a framework designed to select the most informative data point under label scarcity, provides a principled foundation for addressing such challenges (zheng2006selectively; zheng2002active). In particular, active learning allows the analyst to determine which data points should be acquired to maximize model improvement while minimizing redundant or uninformative acquisitions. Active learning strategies are typically divided into two families: optimization-based and heuristic-based approaches (cohn1996active; tong2001active; settles2011theories). Optimization-based methods rely on objective functions that formalize notions of informativeness, such as uncertainty reduction or expected loss. In contrast, heuristic-based approaches draw on approximations or model ensembles to identify uncertain points. In our study, we focus on one strategy from each category: variance-based active learning (VBAL), rooted in optimal experimental design, and Query-by-Committee active learning (QBCAL), a heuristic method based on committee disagreement. As a benchmark, we include a random sampling corrected (RSC) strategy.
@@ 3.3.1 Variance-Based Active Learning (VBAL)
Variance-based active learning (VBAL) is based in Optimal Experimental Design (OED), which provides criteria for selecting data points that most effectively reduce model uncertainty in regression tasks (lu2024data; atkinson1996usefulness; atkinson2001one; atkinson2007optimum). A central quantity in OED is the information matrix $\mathbf{X}^{\top}\mathbf{X}$, derived from the feature $\mathbf{X}$ of the labelled dataset (zheng2006selectively). Several classical criteria operate on this matrix to guide selection: D-optimality seeks to minimize the volume of the confidence ellipsoid around the regression coefficients by maximizing $\det(\mathbf{X}^{\top}\mathbf{X})$; The A-optimality minimizes the average variance of the coefficient estimates by minimizing $\text{tr}(\mathbf{X}^{\top}\mathbf{X})^{-1}$; and the G-optimality focuses on predictive reliability, aiming to minimize the maximum prediction variance in the design space. While D and A-optimality optimize parameter estimation globally, G-optimality is more aligned with our goal: ensuring that predictions are uniformly reliable across all regions of the input space. In formal terms, G-optimality minimizes the worst-case prediction variance, defined as
@@ 3.3.2 Query-by-Committee-Based Active Learning (QBCAL)
Classical QBCAL approaches in classification typically fall into two categories: vote entropy (freund1997selective; argamon1999committee), which quantifies disagreement among classifiers by evaluating the entropy of predicted labels, and margin confidence (abe1998query; settles2011theories), which prioritizes points with the smallest difference between the most probable class labels. Although effective in classification tasks, these strategies are not directly applicable in regression settings where the output is continuous rather than categorical.
@@ 4.1.1 Dataset and case-study
We used a public real estate data set from Sindian District, New Taipei City (yeh2021realestate), containing key transaction features (date, age, MRT distance, convenience stores) and unit-area price. A linear regression model is used to explain the variation in prices.
@@ 4.1.2 What are the effects of applying active learning?
with $d_{0}=0.1$, $d_{1}=0.5$, and $\underline{X_{3}},\overline{X_{3}}]$ the observed range. This serves only to induce heterogeneous seller costs (e.g., privacy, effort, reluctance) in a controlled manner; it does not imply that property characteristics determine economic WTS. Prices follow (5d), with all other parameters identical in BC and SC. Under BC (Figure 3), VBAL and QBCAL reach higher improvement (21.17%, 21.23%) than RSC (20.01%) at similar cost ($\approx\textsterling 10$). Under SC (Figure 4), they again improve faster while costing far less (£4.90–£4.96 vs £10.34). A summary is provided in Table 1.
@@
To confirm that the observed average cost differences are statistically significant, we applied the non-parametric Wilcoxon signed-rank test. The test compared each active learning method (VBAL and QBCAL) with the random sampling corrected (RSC) baseline across 50 paired runs under both pricing schemes. Table 2 reports the two-sided $p$-values and median cost differences. All comparisons yield $p<0.05$, indicating significant differences in cost performance. Under buyer-centric (BC) pricing, VBAL and QBCAL show notably lower median average costs than RSC (16.409-16.409 and 16.499-16.499), while under seller-centric (SC) pricing, both also achieve smaller but consistent reductions (0.059-0.059 and 0.027-0.027). These results confirm that the cost-efficiency advantages of the proposed methods are statistically robust rather than due to random variation.
@@
Strategy Approach $B$ exhausted? $\alpha$ met? Bought Spent (£) Improvement (%) Avg. Cost (£/data)
VBAL BC No Yes 11 10.56 21.17 0.96
SC No Yes 11 4.90 21.17 0.45
QBCAL BC No Yes 11 10.59 21.23 0.96
SC No Yes 11 4.96 21.23 0.45
RSC BC No Yes 20 9.98 20.01 0.50
SC No Yes 20 10.34 20.01 0.52
Table 2: Wilcoxon signed-rank test comparing active learning methods with RSC baseline (estimation-quality-focused scenario)
Approach Comparison $p$-value Median $\Delta$ (Method – RSC)
BC VBAL vs RSC $1.47\times 10^{-8}$ 16.409-16.409
BC QBCAL vs RSC $6.99\times 10^{-9}$ 16.499-16.499
SC VBAL vs RSC $1.35\times 10^{-13}$ 0.059-0.059
SC QBCAL vs RSC $4.01\times 10^{-4}$ 0.027-0.027
Figure 3: Buyer-centric pricing approach (the starting point is not purchasing any data point)
Figure 4: Seller-centric pricing approach (the starting point is not purchasing any data point)
@@ 4.1.3 Does active learning improve cost-efficiency for analysts?
Beyond the averages reported in Table 1, a more revealing picture arises when we track how cost efficiency evolves from the analyst’s perspective. Figure 5 illustrates the cumulative average of the cost-efficiency ratio, $\Delta v/\Delta c$, achieved up to the data points purchased, where $\Delta v$ represents the reduction in posterior variance due to each data point, and $\Delta c$ denotes its associated price. A higher value of $\Delta v/\Delta c$ indicates a greater reduction in variance per unit cost. As shown in this figure, VBAL and QBCAL consistently outperform the baseline RSC method. Note that we report dynamics only for SC because, under constraints (5c)–(5d) in the optimization problem (5), the cost-efficiency ratio in BC is fixed at $1/\phi$ and thus omitted.
Figure 5: Analyst’s side analysis on SC approach: variance reduction per unit cost. Results are represented as cumulative averages (i.e., as the average up to that data point purchased).
@@ 4.1.4 Do pricing approaches affect sellers?
Figure 6: Revenue comparison for sorted data sellers under buyer-centric and seller centric pricing approaches
@@ 4.1.5 Is Our Active Learning Strategy Robust?
A Monte Carlo simulation with 1,000 replications was conducted by repeatedly resampling subsets of the original data, with a model-improvement threshold fixed at 20%. Figures 7 and 8 show the resulting distributions of effectively purchased data points (i.e., data points used to improve the model) for buyer-centric (BC) and seller-centric (SC) pricing schemes. VBAL and QBCAL exhibit concentrated frequency distributions (Figures 77(a)7(b) and 88(a)8(b)), indicating stable behaviour across replications, while the random baseline (RSC) produces flatter distributions (Figures 77(c) and 88(c)), reflecting higher variability and lower robustness.
@@
Parameter Sensitivity.
A second Monte Carlo analysis evaluated how key parameters affect performance in the estimation-quality-focused setting. We varied (i) the WTP of the buyer ($\phi$), (ii) the WTS scaling coefficient of the seller ($d_{1}$), and (iii) the budget limit ($B$), holding all other parameters at baseline (the improvement threshold was raised to 90% to clarify WTP trends). Figures 9(a)9(f) and Table 4 (Appendix) show that VBAL and QBCAL maintain smooth, monotonic purchase patterns, whereas RSC exhibits irregular fluctuations due to its non-adaptive design. As $d_{1}$ increases, stricter acceptance reduces purchases, with occasional local increases when the remaining budget allows high-value purchases. Overall, VBAL and QBCAL acquire fewer but more informative samples than RSC, demonstrating stronger selectivity and cost efficiency.
@@ 4.2.1 Case study and dataset
We use the BDG2 dataset (wang2022adaptive), which contains the hourly energy consumption of 1,636 buildings from 2016 to 2017. For illustration, we consider an educational building (“Rachael”) as the data analyst and identify its most similar building (“Madge”) via Euclidean distance in the feature space, treating it as the data seller. Each observation includes the month, weekday, day, hour, holiday indicator, air temperature, and energy consumption. We adopt a sliding-window approach using only the “value” characteristic (hourly consumption) to construct lagged inputs: the consumption of the past seven days’ at 24-hour intervals (e.g. t168t{-}168, t144t{-}144, …, t24t{-}24), with the current hour’s consumption at time $t$ as the prediction target. Then a linear forecasting model is trained on these lagged features.
@@ 4.2.2 What are the effects of applying active learning?
Table 3 compares the performance of the proposed active learning strategies (VBAL and QBCAL) with the random sampling corrected strategy (RSC). In the BC setting, VBAL and QBCAL achieve similar levels of model improvement ($\sim 17\%$) at lower average cost per data point (£12.81 and £7.68, respectively) relative to RSC (£13.53). In the SC setting, the average cost per data point is identical between methods by design, so performance differences arise solely from the selectivity of the strategies rather than from the pricing. The trends in model improvement and accumulated cost closely mirror those in Section 4.1.2 and are omitted for brevity. Statistical testing for BC is not repeated here as well. Under the seller-centric mechanism, costs are fully determined by fixed willingness-to-sell (WTS) values, making statistical testing not applicable.
@@ 4.2.5 Is our active learning strategy robust?
We ran 1,000 Monte Carlo iterations with a fixed 20% improvement target. Figures 12 and 13 report the distributions of the data points needed under BC and SC, respectively. VBAL and QBCAL show narrow, stable distributions, indicating strong robustness to random data variation, while RSC exhibits substantially wider spread.
@@
We examine robustness through two analyses: (i) one-at-a-time parameter variations (Figure 14) and (ii) a Monte Carlo study over 50 random splits (summary in Table 5, Appendix). Figures 14(a)14(b) show the effect of WTP: in BC, VBAL and QBCAL purchase more data as $\phi$ increases up to 50, then decline once the budget constraint (5e) binds; in SC, effective purchases decrease monotonically with WTP, while RSC fluctuates due to its non-adaptive design. Figures 14(c)14(d) vary the seller’s WTS scale, and all methods buy fewer labels as the WTS increases. Figures 14(e)14(f) vary budget limits (with the improvement threshold raised to 90% to sharpen trends): more budget leads to more acquisitions, with VBAL and QBCAL consistently requiring fewer points than RSC. In general, both VBAL and QBCAL remain selective and cost-effective across all settings.
@@ 5.2 Future works
Efficient data purchase is becoming crucial in our data-driven economies. Our proposed innovative active learning market provides a strong foundation to address this significant problem by offering a trade-off between cost and efficiency. Building on this success, further advancements can enhance its applicability. First, while current data is purchased in a batch setting, real-world data may arrive continuously as a stream, creating opportunities to develop a real-time label market. Second, a valuable extension is to investigate the multiple-buyer-multiple-seller settings which involve competition between sellers and buyers in a resource-constrained environment. In such competitive environments, game-theoretic tools and combinatorial optimization methods (e.g., knapsack heuristics) may play a larger role, complementing active learning in analyzing pricing, market dynamics, and efficiency. Additionally, a one-shot formulation of VBAL is also possible, even in the single-buyer–multiple-seller case, where the budget is allocated in a single step. This connects naturally to knapsack heuristics, but requires the strong assumption that the utilities (or priorities) of all labels are known in advance. For this reason, we focus on iterative allocation, which better reflects adaptive practice, while leaving one-shot variants as an interesting theoretical direction for future work. Third, the design of the market for data and information requires careful reconsideration. Unlike traditional commodities, the uniqueness of data lies in defining both willingness to pay and willingness to sell, as its value is highly context-dependent. For data buyers, their WTP may depend on relevance of tasks, timeliness of sharing, quality of data, among others. For data sellers, their WTS can be influenced by the cost of acquiring data (and storing or sharing it), heterogeneous competition, or privacy constraints. Studying these interplay factors will provide a more comprehensive understanding of how the data market should be structured. Through modelling those dynamics, one can develop more rigorous pricing mechanisms that incorporate strategic behaviour, and thus ensure a fair, efficient, and incentive-compatible data market in the real-world scenarios. Fourth, while our empirical results are strong, we note that deriving formal performance or regret guarantees is particular challenging in our setting. Classical theoretical results in active learning typically assume that all labels have identical unit cost, while our market-based setting allows heterogeneous prices that depend on sellers’ costs or preferences. This means that performance depends jointly on label informativeness and cost, and utilities evolve adaptively as labels are acquired, making standard analyses inapplicable. Developing theoretical guarantees in such heterogeneous-cost settings is therefore a challenging but important direction for future research. Finally, while our framework currently focuses on linear regression and models linear in parameters, extending it to non-convex models such as deep neural networks is an important direction. Estimating label value in such settings is substantially more challenging due to non-convex optimization landscapes, stochastic training procedures, and state dependence. Possible avenues include using influence functions, gradient-based embeddings, or Bayesian uncertainty estimates to approximate label utilities.

### v2

@@ 3.3 Data purchase strategies
The data purchase problem addressed in this paper is inherently constrained by budget and governed by heterogeneous pricing. Active learning, a framework designed to select the most informative data point under label scarcity, provides a principled foundation for addressing such challenges (Zheng and Padmanabhan, 2006, 2002). In particular, active learning allows the analyst to determine which data points should be acquired to maximize model improvement while minimizing redundant or uninformative acquisitions. Active learning strategies are typically divided into two families: optimization-based and heuristic-based approaches (Cohn et al., 1996; Tong, 2001; Settles, 2011). Optimization-based methods rely on objective functions that formalize notions of informativeness, such as uncertainty reduction or expected loss. In contrast, heuristic-based approaches draw on approximations or model ensembles to identify uncertain points. In our study, we focus on one strategy from each category: variance-based active learning (VBAL), rooted in optimal experimental design, and Query-by-Committee active learning (QBCAL), a heuristic method based on committee disagreement. As benchmarks, we include RSC and a Greedy Knapsack (GK) baseline that greedily purchases the lowest-cost candidates by ranking them using $1/\eta_{j}$, since $l_{j}$ is not observable ex ante.
@@ 3.3.1 Variance-Based Active Learning (VBAL)
Variance-based active learning (VBAL) is based in Optimal Experimental Design (OED), which provides criteria for selecting data points that most effectively reduce model uncertainty in regression tasks (Lu et al., 2024; Atkinson, 1996; Atkinson and Bailey, 2001; Atkinson et al., 2007). A central quantity in OED is the information matrix $\mathbf{X}^{\top}\mathbf{X}$, derived from the feature $\mathbf{X}$ of the labelled dataset (Zheng and Padmanabhan, 2006). Several classical criteria operate on this matrix to guide selection: D-optimality seeks to minimize the volume of the confidence ellipsoid around the regression coefficients by maximizing $\det(\mathbf{X}^{\top}\mathbf{X})$; The A-optimality minimizes the average variance of the coefficient estimates by minimizing $\text{tr}(\mathbf{X}^{\top}\mathbf{X})^{-1}$; and the G-optimality focuses on predictive reliability, aiming to minimize the maximum prediction variance in the design space. While D and A-optimality optimize parameter estimation globally, G-optimality is more aligned with our goal: ensuring that predictions are uniformly reliable across all regions of the input space. In formal terms, G-optimality minimizes the worst-case prediction variance, defined as
@@ 3.3.2 Query-by-Committee-Based Active Learning (QBCAL)
Classical QBCAL approaches in classification typically fall into two categories: vote entropy (Freund et al., 1997; Argamon-Engelson and Dagan, 1999), which quantifies disagreement among classifiers by evaluating the entropy of predicted labels, and margin confidence (Abe and Mamitsuka, 1998; Settles, 2011), which prioritizes points with the smallest difference between the most probable class labels. Although effective in classification tasks, these strategies are not directly applicable in regression settings where the output is continuous rather than categorical.
@@ 4.1.1 Dataset and case-study
We used a public real estate data set from Sindian District, New Taipei City (Yeh, 2021), containing key transaction features (date, age, MRT distance, convenience stores) and unit-area price. A linear regression model is used to explain the variation in prices.
@@ 4.1.2 What are the effects of applying active learning?
with $d_{0}=0.1$, $d_{1}=0.5$, and $\underline{X_{3}},\overline{X_{3}}]$ the observed range. This serves only to induce heterogeneous seller costs (e.g., privacy, effort, reluctance) in a controlled manner; it does not imply that property characteristics determine economic WTS. Prices follow (5d), with all other parameters identical in BC and SC. Under BC (Figure 3(b)), VBAL reaches higher improvement (21.17%) than RSC (20.01%) and GK (20.67%) at similar cost ($\approx\textsterling 10$). Under SC (Figure 3(c)), both VBAL and QBCAL improve faster while costing far less (£4.56–£3.89 vs £10.08-£7.96). A summary is provided in Table 1.
@@
To assess whether the differences observed in Table 1 are statistically significant, we applied the non-parametric Wilcoxon signed-rank test on improvement per purchased point (%/point), with significance level $p=0.05$. The test compares paired outcomes across 50 random splits for each method under both buyer-centric (BC) and seller-centric (SC) pricing. Table 2 reports Wilcoxon test results for improvement per purchased point. Under BC pricing, VBAL significantly outperforms the RSC baseline, while QBCAL vs RSC is not statistically significant. Against GK, both VBAL and QBCAL achieve significantly higher per-point improvement under both BC and SC pricing, indicating that GK’s low-price selection does not translate into stronger improvement per purchased point.
@@
Strategy Approach $B$ exhausted? $\alpha$ met? Purchased Spent (£) Improvement (%) Avg. Impro. (%/point)
VBAL BC No Yes 11 10.56 21.17 1.92
SC 4.56
QBCAL BC No Yes 10 10.08 20.22 2.02
SC 3.89
RSC BC No Yes 20 9.98 20.01 1.00
SC 10.08
GK BC No Yes 30 10.31 20.67 0.69
SC 7.96
Table 2: Wilcoxon signed-rank test comparing improvement per purchased point (estimation-quality-focused scenario)
Approach Comparison $p$-value Median $\Delta$ (%/point)
BC VBAL vs RSC $3.349\times 10^{-8}$ $0.7748$
BC QBCAL vs RSC $1.28\times 10^{-1}$ $0.1988$
BC VBAL vs GK $9.521\times 10^{-13}$ $1.0428$
BC QBCAL vs GK $1.045\times 10^{-5}$ $0.4370$
SC VBAL vs RSC $3.706\times 10^{-9}$ $0.7748$
SC QBCAL vs RSC $5.547\times 10^{-2}$ $0.1988$
SC VBAL vs GK $5.862\times 10^{-14}$ $1.0428$
SC QBCAL vs GK $9.364\times 10^{-7}$ $0.4370$
Figure 3: Different pricing approach (the starting point is not purchasing any data point)
@@ 4.1.3 Does active learning improve cost-efficiency for analysts?
Beyond the average improvements reported in Table 1, a more revealing picture arises when we track how cost efficiency evolves from the analyst’s perspective. Figure 4 illustrates the cumulative average of the cost-efficiency ratio, $\Delta v/\Delta c$, achieved up to the data points purchased, where $\Delta v$ represents the reduction in posterior variance due to each data point, and $\Delta c$ denotes its associated price. A higher value of $\Delta v/\Delta c$ indicates a greater reduction in variance per unit cost. As shown in this figure, VBAL and QBCAL consistently outperform the baseline RSC method. Greedy Knapsack (GK) exhibits a high initial cumulative $\Delta v/\Delta c$ mainly because it purchases the lowest-cost points first ($1/\eta_{j}$ in a descending order), which inflates the early running average. However, this advantage quickly diminishes as additional low-cost points yield smaller marginal variance reductions, causing GK’s cumulative $\Delta v/\Delta c$ to decline. Consistently, Table 1 shows that GK delivers lower improvement per purchased point overall than VBAL and QBCAL. Note that we report dynamics only for SC because, under constraints (5c)–(5d) in the optimization problem (5), the cost-efficiency ratio in BC is fixed at $1/\phi$ and thus omitted
Figure 4: Analyst’s side analysis on SC approach: variance reduction per unit cost. Results are represented as cumulative averages (i.e., as the average up to that data point purchased)
@@ 4.1.4 Do pricing approaches affect sellers?
Figure 5: Revenue comparison for sorted data sellers under BC and SC pricing approaches
@@ 4.1.5 Is Our Active Learning Strategy Robust?
A Monte Carlo simulation with 1,000 replications was conducted by repeatedly resampling subsets of the original data, with a model-improvement threshold fixed at 20%. Figures 6 and 7 show the resulting distributions of effectively purchased data points (i.e., data points used to improve the model) for buyer-centric (BC) and seller-centric (SC) pricing schemes. VBAL and QBCAL exhibit concentrated frequency distributions (Figures 66(a)6(b) and 77(a)7(b)), indicating stable behaviour across replications, while the random (RSC) and the Greedy Knapsack (GK) baselines produce flatter distributions (Figures 66(c)6(d) and 77(c)7(d)), reflecting higher variability and lower robustness.
Parameter Sensitivity.
We further examine parameter sensitivity in the estimation-quality-focused setting by varying (i) the buyer’s WTP $\phi$, (ii) the seller’s WTS scaling coefficient $d_{1}$, and (iii) the budget limit $B$, while keeping all other parameters at their baseline values. Figures 8(a)8(f) and Table 4 (Appendix) show that the effective number of purchased points is largely insensitive to $\phi$ in the tested range under both BC and SC pricing: VBAL and QBCAL consistently purchase only a small set of informative points (around 10), whereas RSC and Greedy Knapsack purchase substantially more points (around 19–21 and $28$, respectively). Increasing $B$ from £10 to £30 has a limited impact in these experiments, as most methods reach the variance threshold before exhausting the larger budgets.
@@ 4.2.1 Case study and dataset
We use the BDG2 dataset (Wang et al., 2022), which contains the hourly energy consumption of 1,636 buildings from 2016 to 2017. For illustration, we consider an educational building (“Rachael”) as the data analyst and identify its most similar building (“Madge”) via Euclidean distance in the feature space, treating it as the data seller. Each observation includes the month, weekday, day, hour, holiday indicator, air temperature, and energy consumption. We adopt a sliding-window approach using only the “value” characteristic (hourly consumption) to construct lagged inputs: the consumption of the past seven days’ at 24-hour intervals (e.g. t168t{-}168, t144t{-}144, …, t24t{-}24), with the current hour’s consumption at time $t$ as the prediction target. Then a linear forecasting model is trained on these lagged features.
@@ 4.2.2 What are the effects of applying active learning?
Table 3 compares the performance of the proposed active learning strategies (VBAL and QBCAL) with the random sampling corrected strategy (RSC). In the BC setting, VBAL and QBCAL achieve similar levels of model improvement ($\sim 17\%$) at lower average cost per data point (£12.81 and £7.68, respectively) relative to RSC (£13.53). In the SC setting, the average cost per data point is identical between methods by design, so performance differences arise solely from the selectivity of the strategies rather than from the pricing. The trends in model improvement and accumulated cost closely mirror those in Section 4.1.2 and are omitted for brevity. Under the seller-centric mechanism the reported average cost is determined by the fixed WTS schedule, so a Wilcoxon signed-rank test on cost differences is uninformative. For the same reason, we do not include a Greedy Knapsack baseline for this dataset as it cannot meaningfully differentiate points (any selection is equivalent).
@@ 4.2.5 Is our active learning strategy robust?
We ran 1,000 Monte Carlo iterations with a fixed 20% improvement target. Figures 11 and 12 report the distributions of the data points needed under BC and SC, respectively. VBAL and QBCAL show narrow, stable distributions, indicating strong robustness to random data variation, while RSC exhibits substantially wider spread.
@@
We examine robustness through two analyses: (i) one-at-a-time parameter variations (Figure 13) and (ii) a Monte Carlo study over 50 random splits (summary in Table 5, Appendix). Figures 13(a)13(b) show the effect of WTP: in BC, VBAL and QBCAL purchase more data as $\phi$ increases up to 50, then decline once the budget constraint (5e) binds; in SC, effective purchases decrease monotonically with WTP, while RSC fluctuates due to its non-adaptive design. Figures 13(c)13(d) vary the seller’s WTS scale, and all methods buy fewer labels as the WTS increases. Figures 13(e)13(f) vary budget limits (with the improvement threshold raised to 90% to sharpen trends): more budget leads to more acquisitions, with VBAL and QBCAL consistently requiring fewer points than RSC. In general, both VBAL and QBCAL remain selective and cost-effective across all settings.
@@ 5.2 Future works
Efficient data purchase is becoming crucial in our data-driven economies. Our proposed innovative active learning market provides a strong foundation to address this significant problem by offering a trade-off between cost and efficiency. Building on this success, further advancements can enhance its applicability. First, while current data is purchased in a batch setting, real-world data may arrive continuously as a stream, creating opportunities to develop a real-time label market. Second, a valuable extension is to investigate the multiple-buyer-multiple-seller settings which involve competition between sellers and buyers in a resource-constrained environment. In such competitive environments, game-theoretic tools and combinatorial optimization methods (e.g., knapsack heuristics) may play a larger role, complementing active learning in analyzing pricing, market dynamics, and efficiency. Additionally, a one-shot formulation of VBAL is also possible, even in the single-buyer–multiple-seller case, where the budget is allocated in a single step. This connects naturally to knapsack heuristics, but requires the strong assumption that the utilities (or priorities) of all labels are known in advance. For this reason, we focus on iterative allocation, which better reflects adaptive practice, while leaving one-shot variants as an interesting theoretical direction for future work. Third, the design of the market for data and information requires careful reconsideration. Unlike traditional commodities, the uniqueness of data lies in defining both willingness to pay and willingness to sell, as its value is highly context-dependent. For data buyers, their WTP may depend on relevance of tasks, timeliness of sharing, quality of data, among others. For data sellers, their WTS can be influenced by the cost of acquiring data (and storing or sharing it), heterogeneous competition, or privacy constraints. Studying these interplay factors will provide a more comprehensive understanding of how the data market should be structured. Through modelling those dynamics, one can develop more rigorous pricing mechanisms that incorporate strategic behaviour, and thus ensure a fair, efficient, and incentive-compatible data market in the real-world scenarios. Fourth, while our empirical results are strong, we note that deriving formal performance or regret guarantees is particularly challenging in our setting. Classical theoretical results in active learning typically assume that all labels have identical unit cost, while our market-based setting allows heterogeneous prices that depend on sellers’ costs or preferences. This means that performance depends jointly on label informativeness and cost, and utilities evolve adaptively as labels are acquired, making standard analyses inapplicable. Developing theoretical guarantees in such heterogeneous-cost settings is therefore a challenging but important direction for future research. More concrete partial insights may still be possible under stylised assumptions, e.g., bounded or i.i.d. cost heterogeneity, and informativeness proxies that are bounded and (approximately) unbiased for the true marginal improvement. In such regimes, one could aim for comparison-based guarantees against random baselines, or one-step (myopic) optimality/Pareto-efficiency properties of cost-weighted greedy rules. Finally, while our framework currently focuses on linear regression and models linear in parameters, extending it to non-convex models such as deep neural networks is an important direction. Estimating label value in such settings is substantially more challenging due to non-convex optimization landscapes, stochastic training procedures, and state dependence. Possible avenues include using influence functions, gradient-based embeddings, or Bayesian uncertainty estimates to approximate label utilities.